The 600 students at King Middle School are divided into 3 groups of equal size for lunch. Each group has lunch at a different time. A computer randomly assigns each student to one of 3 lunch groups. Find the probability that 3 friends, AL,Bob,Carol will be asigned to the same lunch group.
Found 2 solutions by stanbon, longjonsilver:
Answer by stanbon(75887)

1/[600C3]=1/35,820,200
Cheers,
Stan H.

Answer by longjonsilver(2297)

P(student in a particular group) = 1/3

So, P(all 3 in first group) = (1/3)*(1/3)*(1/3) = 1/27
Also, P(all 3 in second group) = 1/27 too
and P(all 3 in third group) = 1/27 also

So, P(3 friends are all together) = 1/27 + 1/27 + 1/27
P(3 friends are all together) = 3/27
P(3 friends are all together) = 1/9

jon.